4x+6y=12
2x+3y=6
what is the relationship, if any, between 12 and 6?

Answer by ewatrrr(24785)

 
Hi
4x+6y=12 and 2x+3y=6
NOte: 4x+6y=12 is 2(2x+3y)= 2•6